Let f(x) = 1 /2 ,0< x< 1 or 2< x< 3 ,zero elsewhere,be the pdf

of X. Find \pi _{0.70} . Show your work!
Answer #1

here let a is the 70th percentile ;

hence \int_{0}^{1} f(x) dx +\int_{2}^{a} f(x) dx =0.7

\int_{0}^{1}(1/2) dx +\int_{2}^{a}(1/2) dx =0.7

(x/2)|10 +(x/2)|a2 =0.7

0.5+(a-2)/2 =0.7

a-2=2*(0.7-0.5)

a=2+0.4

a=\pi0.70=2.4